### v1

# MonoDGP: Monocular 3D Object Detection with Decoupled-Query and Geometry-Error Priors

Abstract
Perspective projection has been extensively utilized in monocular 3D object detection methods. It introduces geometric priors from 2D bounding boxes and 3D object dimensions to reduce the uncertainty of depth estimation. However, due to depth errors originating from the object’s visual surface, the height of the bounding box often fails to represent the actual projected central height, which undermines the effectiveness of geometric depth. Direct prediction for the projected height unavoidably results in a loss of 2D priors, while multi-depth prediction with complex branches does not fully leverage geometric depth. This paper presents a Transformer-based monocular 3D object detection method called MonoDGP, which adopts perspective-invariant geometry errors to modify the projection formula. We also try to systematically discuss and explain the mechanisms and efficacy behind geometry errors, which serve as a simple but effective alternative to multi-depth prediction. Additionally, MonoDGP decouples the depth-guided decoder and constructs a 2D decoder only dependent on visual features, providing 2D priors and initializing object queries without the disturbance of 3D detection. To further optimize and fine-tune input tokens of the transformer decoder, we also introduce a Region Segment Head (RSH) that generates enhanced features and segment embeddings. Our monocular method demonstrates state-of-the-art performance on the KITTI benchmark without extra data. Code is available at https://github.com/PuFanqi23/MonoDGP.
1 Introduction
Figure 1: Comparison with MonoDETR. (b) MonoDGP proposes the RSH for enhanced features and segment embeddings, along with an independent visual decoder for 2D detection and query initialization. (c) Furthermore, we introduce a geometry error prior to convert uneven depth distribution into a more concentrated depth error distribution.
3D object detection has always been a critical research topic in the fields of autonomous driving and robotic perception. Compared with LiDAR- [17, 35, 36, 47] or stereo-based [19, 38, 30, 18, 9] methods, monocular 3D object detection exhibits lower spatial positional accuracy without depth information. Owing to their cost-effectiveness and ease of configuration [21, 31], monocular methods have remained a primary focus of research in recent years.
Without extra data, the direct prediction of 3D objects from single 2D images is an extremely challenging problem. During the formation of 2D images, crucial depth-related information is compressed, resulting in an inherently ill-posed issue for direct depth estimation [32, 53, 28, 50, 21]. This limitation significantly undermines the accuracy of 3D object detection. However, under certain conditions, the uncertainty of depth estimation can be reduced with substantial prior knowledge. Particularly in the realm of autonomous driving, where the major objects of interest, such as cars, are rigid entities with known geometric properties.
Previous methods [27, 44, 26, 49, 53] typically calculate geometric depth through the projection formula based on bounding boxes and the dimension heights. However, since 2D images are inaccurate in capturing depth errors across an object’s visual surface, the bounding box height is often larger than the projected central height. To address the limitations inherent in single-depth prediction, some approaches[49, 21, 48] integrate multi-depth predictions, including direct depth and geometric depth, and generate final results through weighted averaging. Nevertheless, MonoCD [48] indicates that if the multiple predicted depths cluster on the same side of the ground truth depth, even weighted averaging fails to correct the depth error. DID-M3D [28] decomposes raw depth into visual depth and attribute depth, but ignores geometric depth, using perspective transformation only for data augmentation. Similarly, the geometric depth can be regarded as the visual surface depth, whose value depends on the object’s appearance and position [10]. While the depth error is determined by the inherent attributes of the object and remains invariant with changes in perspective. Besides, the distribution of depth error is more concentrated compared to raw depth. Utilizing error prediction instead of direct prediction is a simple and effective improvement. This approach can not only alleviate the long-tail effect of raw depth distribution, but also avoid the problem where multiple depths fail to complement each other after weighting.
After introducing the geometric prior, we observed that the accuracy of the bounding boxes is often poor during the initial training stage. This may lead the model to associate depth with incorrect bounding boxes, thereby complicating the convergence process. To address this issue, we decouple the depth-guided decoder proposed by MonoDETR [49] and construct a 2D visual decoder for independent 2D detection and query initialization. Additionally, we propose a Region Segmentation Head (RSH) to strengthen the foreground features and provide segment embeddings for semantically similar input tokens. By emulating BERT [14]’s approach to handling distinct components of textual inputs, we treat the target region as one “sentence” and the background region as another “sentence”.
In summary, we propose a transformer-based monocular detector with decoupled queries and geometric errors, called MonoDGP. Our contributions are listed as follows:
  • We adopt perspective-invariant error prediction to replace multi-depth prediction, effectively alleviating the long-tail effect of depth distribution and reducing learning complexities.
  • We construct a decoupled visual decoder to initialize queries and reference points of the depth-guided decoder, boosting the model convergence and training stability.
  • We present a Region Segmentation Head (RSH) to enhance object features and suppress background noise. Additionally, segment embeddings are incorporated during depth encoding to refine contextual representations.
  • Experiments illustrate the superiority of MonoDGP. Evaluated on the KITTI 3D object detection benchmark, without any extra data, our method achieves the state-of-the-art (SOTA) performance among monocular detectors.
Figure 2: Overall architecture of our MonoDGP. The network comprises three components: feature extraction and enhancement, transformer encoder-decoder and object detection heads. The parallel visual and depth branches within the transformer are represented by yellow and blue, respectively. The predicted depth is separated into geometric depth $Z_{geo}$, obtained from the perspective projection formula, and depth error $Z_{err}$ to correct the formula.
2 Related work
Extra training data.  The main challenge of monocular camera-based methods is insufficient of depth cues, which renders 3D object detection an ill-posed problem. Consequently, some approaches [42, 33, 20, 13, 28, 45] incorporate additional data, such as point clouds or depth maps, enabling the model to implicitly learn depth features during training. MonoPGC [45] introduces pixel-level depth map prediction as an auxiliary task, encoding 3D coordinates into depth-perception features to perceive the spatial positions of pixels. CaDDN [33] estimates the classified depth distribution for each pixel, generates high-quality bird’s-eye view feature representations, and achieves accurate 3D object predictions. OccupancyM3D [29] creates occupancy labels from point clouds, and monitors occupancy status in frustum space and 3D space. Given the reliance on expensive sensors and complex collection processes for supplementary information, MonoDGP exclusively utilizes monocular images during the training and inference stages, thereby enhancing the model’s practicality and deployment feasibility.
Transformer-based methods.  Previous works [50, 37, 26, 28, 23, 21] have employed convolutional neural networks (CNNs) to extract local, fine-grained information at various spatial levels. Although CNN-based methods demonstrate superior speed and efficiency, they require multi-layer convolution to progressively aggregate features for capturing long-distance dependencies. Additionally, non-maximum suppression (NMS) technology is employed to filter out redundant detection boxes. Benefiting from the global perception and dynamic adjustment abilities of the attention mechanism [41], an increasing number of studies [13, 49, 52] have shifted their focus toward Transformer-based monocular detectors. For instance, MonoDTR [13] proposes to conduct depth position encoding and integrates global depth information into the Transformer to guide the detection. MonoDETR [49] adheres to the DETR [5] architecture, utilizing object-level depth labels to predict the foreground depth map, constructing a depth-guided decoder, and incorporating object queries for global feature aggregation. MonoDGP also adopts the DETR framework, initializing queries with a decoupled 2D visual decoder and introducing segment embeddings to enhance the contextual relationships among pixels.
Multi-depth estimation modes.  Early monocular methods [8, 39, 3, 22, 24] directly estimate object depth through end-to-end network training. Many researchers [1, 2, 4, 51, 15, 37, 43] extend the depth estimation branches by introducing perspective projection and mathematical priors. For example, GUPNet [26] proposes a geometry uncertainty projection module and employs a Laplacian distribution to represent geometric depth with correlation-learned bias. To mitigate uncertainties in single depth estimation methods, MonoDDE [21] introduces key points into 20 different depth branches, emphasizing the significance of diversity in depth prediction. MonoDETR [49] merges predicted results from direct depth, object-level depth map, and geometric depth. MonoCD [48] further discusses the complementarity among multiple depths and combines them to cancel out positive or negative errors. However, the geometric depth and errors (such as depth/height error) in the projection formula have not been fully explored. Based on a comprehensive analysis of geometric priors, MonoDGP decouples raw depth to obtain perspective-invariant depth error, providing a superior alternative to multi-depth approaches.
3 Method
Figure 3: The structure of region segmentation module. The multi-scale feature maps are progressively upsampled and added, outputting target region probabilities by segment heads. Subsequently, enhanced features are achieved through element-wise multiplication with original features. Finally, segment embeddings are acquired under threshold constraints.
3.1 Overview
The overall framework of MonoDGP is shown in Fig. 2. Firstly, a single-view image is processed by ResNet50 to extract multi-scale feature maps ${f_{i}}\in{\mathbb{R}^{\frac{H}{{{2^{i+3}}}}\times\frac{W}{{{2^{i+3}}}}\times C}},i=0,1,2,3$. The RSH module then predicts target region probabilities and enhances visual features $f_{V}$. A lightweight depth predictor subsequently estimates the depth map and generates depth features $f_{D}$. Both the enhanced visual features $f_{V}$ and depth features $f_{D}$ are fed into two parallel transformer encoders. The transformer decoder consists of a 2D visual decoder and a 3D depth-guided decoder. In the 2D decoder, object queries adaptively achieve semantic scene understanding from visual embeddings, while the 3D decoder extracts geometric and spatial relationships from both visual and depth embeddings. Ultimately, we utilize decoupled queries as inputs for 2D/3D detection heads separately. It is worth noting that after calculating geometric depth based on the projection formula, a geometry error correction mechanism is applied to improve the accuracy of depth estimation.
3.2 Region Segmentation
The Region Segmentation Head (RSH) is specifically designed to enhance the fine-grained foreground feature extraction. As shown in Fig. 3, the RSH follows the structure of U-Net [34] and predicts the target probability for each pixel. It is essential for the depth predictor to concentrate more on the perspective and geometric cues within the foreground area. Therefore, the RSH provides enhanced target features and segment embeddings for depth encoding. Technical details are presented below:
Region Segmentation and Enhancement.  We initially utilize a ResNet backbone to downsample and extract multi-scale visual features $f_{0}$ to $f_{3}$. Subsequently, We restore $f_{3}$ to high-resolution feature maps through upsampling operations and skip connections. This integration of shallow and deep features improves multi-scale representation. Specifically, we combine the upsampled $f_{3}$ with $f_{2}$ after a $1\times 1$ convolution and GroupNorm layer. This process is repeated to obtain fused feature maps at various scales. We then employ an SEBlock for attention weighting on each scale’s feature maps. The SEBlock implements a channel attention mechanism through two $1\times 1$ convolution layers: the first convolution reduces the number of feature channels to $1/16$, while the second convolution restores the original number of channels. Finally, we achieve several semantic segmentation maps with the sigmoid function. By multiplying the predicted target probability with original features, RSH effectively enhances foreground features and suppresses background noise.
Figure 4: (a) Schematic illustration of how geometry error occurs. Due to the height of the bounding box $h_{\text{bbox}}$ is larger than the projected central height $h_{c}$, there will be an error between the geometric depth and the ground truth depth. (b) Vehicles at different angles in a bird’s-eye view. The geometric depth $Z_{\text{geo}}$is the distance between the camera plane and the parallel plane on which the closest wheel sits, while the depth error $Z_{\text{err}}$ only depends on dimensions and orientation of the object.
Region loss.  We construct target region masks and set the areas within bounding boxes to 1. Guided by these region masks, we calculate the Dice loss for the target region probabilities at each scale and weight the sum of these losses. The Dice loss, which is based on set similarity, addresses class imbalance and improves segmentation precision. Its definition is as follows:
$L_{region}^{i}=1-\frac{{2\cdot\sum\limits_{j=1}^{{N_{i}}}{{p_{j}}\cdot{g_{j}}}}}{{\sum\limits_{j=1}^{{N_{i}}}{{p_{j}}}+\sum\limits_{j=1}^{{N_{i}}}{{g_{j}}}}}$ (1)
Where $p_{j}$ denotes the predicted probability of the $j$-th pixel, $g_{j}$ represents the $j$-th pixel in the ground truth label, and $N_{i}$ indicates the total number of pixels in the $i$-th feature map ${f_{i}},i=0,1,2,3$. By minimizing the Dice loss, we can directly optimize the overlapping areas between the predicted probabilities and the region masks, enhancing both segmentation accuracy and consistency.
Segment embeddings.  The RSH module generates segment embeddings from target region probabilities that exceed a threshold of 0.5. We subsequently add these segment embeddings to the token and position embeddings. This approach is inspired by the segment embeddings mechanism used in the BERT model. Segment embeddings help to better distinguish the semantic relationships among image tokens, similar to how different “sentences” are represented in BERT.
3.3 Decoupled Query
After obtaining visual embeddings $f_{V}^{e}$ and depth embeddings $f_{D}^{e}$ from two parallel encoders, MonoDGP decouples the depth-guided decoder into a 2D decoder and a 3D decoder. By initializing queries and reference points with the 2D visual decoder, the object queries can acquire spatial position, shape and category information, reducing depth uncertainties from geometric priors. These 2D queries are then fed into the 3D decoder to interact with visual and depth features, allowing each decoder to focus on its own specific task and ensuring stable convergence during training.
2D detection decoder.  We take a set of learnable object queries $q\in{\mathbb{R}^{N\times C}}$ as input, which sequentially passes through an inter-query self-attention layer, a multi-scale deformable cross-attention layer [54], and a feedforward network(FFN), where $N$ denotes the pre-defined maximum number of objects in a single image. The 2D queries ${q_{2d}}$ interact with visual embeddings $f_{V}^{e}$ to extract 2D image features, including color, edges, and texture. The queries and reference points provide an initial perception of spatial positioning. The 3D decoder can utilize these 2D query priors to more rapidly focus on the target area in complex scenes.
3D detection decoder.  The initialized 2D object queries ${q_{2d}}\in{R^{N\times C}}$ are input to the 3D depth-guided decoder. Each decoder block contains a depth cross-attention layer, an inter-query self-attention layer, a visual cross-attention layer and an FFN. Specifically, object queries first extract geometric and perspective cues from depth embeddings $f_{D}^{e}$. Subsequently, the depth-aware queries are sent to the inter-query self-attention layer to promote information exchange. After the visual cross-attention layer and FFN, we generate final 3D queries ${q_{3d}}$ that combine semantic and geometric information, obtaining a more accurate scene-level spatial understanding.
3.4 Geometry Error
Direct depth estimation is often ill-posed in monocular methods. We apply the perspective projection formula $Z=f*H_{3D}/h_{2D}$ to introduce geometric priors and mitigate depth uncertainty. When utilizing the projection formula, it is essential to make sure that the object’s dimension height (such as $AB$ and $CD$ in Fig. 4(a)) is parallel to the projection plane, namely the X-Y plane in the camera coordinate system, and remains perpendicular to the X-Z plane. To guarantee the validity of the formula, we establish the following assumptions:
  • The roll and pitch angles of the monocular camera are set to 0 degrees by default.
  • The influence of ground slope can be overlooked when calculating geometric depth.
Under the given assumptions, the projected central height $h_{c}$ is solely dependent on the 3D height and depth, while the height of the 2D bounding box $h_{\text{bbox}}$ can be influenced by additional factors, including its shape and surface characteristics. As illustrated in Fig. 4 (a), the visual disparity typically causes a depth error $Z_{\text{err}}$ between the actual depth $Z$ and the geometric depth $Z_{\text{geo}}$, which can be formulated as:
$Z=\frac{{f\cdot{Y_{AB}}}}{{{v_{AB}}}}=\frac{{f\cdot H}}{{{h_{c}}}},{Z_{geo}}=\frac{{f\cdot{Y_{CD}}}}{{{v_{CD}}}}=\frac{{f\cdot H}}{{{h_{bbox}}}}$ (2)
$Z={Z_{geo}}+{Z_{err}}$ (3)
where f denotes the focal length of the camera.
The $Z_{\text{geo}}$ reflects the visible surface depth, while $Z_{\text{err}}$ is only related to the inherent attributes of the object. In comparison to the raw depth $Z$, $Z_{\text{err}}$ is perspective-invariant with less spatial uncertainty. We replace the direct depth prediction with the geometry error prediction, and our depth loss is calculated as follows:
${L_{depth}}=\frac{{\sqrt{2}}}{{{\sigma_{d}}}}{\left\|{{Z_{geo}}+{Z_{err}}-{Z_{gt}}}\right\|_{1}}+\log({\sigma_{d}})$ (4)
The perspective projection formula contains two variables: the height of the 2D bounding box and the height of the 3D size. To correct the formula, the depth error is added to the geometric depth for the final prediction, while the errors from both heights can also compensate for the geometric depth. Their error predictions are formulated as:
$Z=\frac{{f\cdot(H+{H_{err}})}}{{{h_{bbox}}}}={Z_{geo}}+\frac{f}{{{h_{bbox}}}}\cdot{H_{err}}$ (5)
$Z=\frac{{f\cdot H}}{{{h_{c}}}}=\frac{{f\cdot H}}{{{h_{bbox}}-{h_{err}}}}$ (6)
These errors can significantly improve the performance of geometric depth.
Figure 5: Comparison of 3D data distributions. With the total number of objects remaining constant, each distribution is visualized using a histogram.
Table 1: Comparisons with state-of-the-art monocular methods on the KITTI test and val sets for the car category. We bold the best results and underline the second-best results. The blue refers to the gain and the red is the decrease of our method.
Test, $AP_{BEV}$ Test, $AP_{3D}$ Val, $AP_{BEV}$ Val, $AP_{3D}$
Method Extra data Reference Easy Mod. Hard Easy Mod. Hard Easy Mod. Hard Easy Mod. Hard
CaDDN [33] CVPR 2021 27.94 18.91 17.19 19.17 13.41 11.46 - - - 23.57 16.31 13.84
MonoDTR [13] CVPR 2022 28.59 20.38 17.14 21.99 15.39 12.73 33.33 25.35 21.68 24.52 18.57 15.51
DID-M3D [28] ECCV 2022 32.95 22.76 19.83 24.40 16.29 13.75 31.10 22.76 19.50 22.98 16.12 14.03
OccupancyM3D [29] LiDAR CVPR 2024 35.38 24.18 21.37 25.55 17.02 14.79 35.72 26.60 23.68 26.87 19.96 17.15
MonoPGC [45] ICRA 2023 32.50 23.14 20.30 24.68 17.17 14.14 34.06 24.26 20.78 25.67 18.63 15.65
OPA-3D [40] Depth RAL 2023 33.54 22.53 19.22 24.60 17.05 14.25 33.80 25.51 22.13 24.97 19.40 16.59
GUPNet [26] ICCV 2021 - - - 20.11 14.20 11.77 31.07 22.94 19.75 22.76 16.46 13.72
MonoCon [23] AAAI 2022 31.12 22.10 19.00 22.50 16.46 13.95 - - - 26.33 19.01 15.98
DEVIANT [16] ECCV 2022 29.65 20.44 17.43 21.88 14.46 11.89 32.60 23.04 19.99 24.63 16.54 14.52
MonoDETR [49] ICCV 2023 33.60 22.11 18.60 25.00 16.47 13.58 37.86 26.95 22.80 28.84 20.61 16.38
MonoCD [48] CVPR 2024 33.41 22.81 19.57 25.53 16.59 14.53 34.60 24.96 21.51 26.45 19.37 16.38
FD3D [46] None AAAI 2024 34.20 23.72 20.76 25.38 17.12 14.50 36.98 26.77 23.16 28.22 20.23 17.04
Ours None - 35.24 25.23 22.02 26.35 18.72 15.97 39.40 28.20 24.42 30.76 22.34 19.02
Improvement - - -0.14 +1.05 +0.65 +0.80 +1.55 +1.18 +1.54 +1.25 +1.33 +1.92 +1.73 +1.87
3.5 Why Geometry Error Works?
The geometry error eliminates perspective correlation of raw depth distribution, thus reducing the learning difficulty of the network. Specifically, the reasons why geometry error works can be analyzed from three aspects:
  • 2D and 3D attributes: why we utilize $h_{bbox}$ instead of $h_{c}$ in the projection formula?
  • Different 3D attributes: why raw depth preforms worse than other 3D attributes?
  • Different geometry errors: why we select depth error?
2D and 3D attributes.  The generation of bounding boxes relies on 2D visual features like color, edges, textures, and shapes. In contrast, predicting 3D attributes requires extra factors including depth positioning and geometric perspective, which are difficult to obtain from single-view images. Benefiting from the network’s powerful capability of marginal feature extraction, $h_{bbox}$ can provide more accurate 2D priors. However, we can hardly predict $h_{c}$ from direct marginal features. The geometric depth calculated by $h_{bbox}$ should be regarded as a reliable surface depth, rather than an inaccurate central depth.
Figure 6: The boxplot of depth-related attributes. To conduct unified comparisons, the data distributions of different attributes are standardized by $X^{\prime}=(X-\mu)/\sigma$ in the boxplot.
Different 3D attributes.  In the absence of direct depth information, the 3D data distribution will determine the learning complexity of the network. The Fig. 5 and 6 show the data distributions and boxplot of depth-related attributes. Based on these figures, we can further evaluate and compare the learning difficulty across different data distributions. The central depth usually exhibits a long-tailed distribution characterized by a wide and uneven range. Conversely, the 3D size distribution tends to be more concentrated. For instance, most cars are around 1.5 meters tall, which facilitates the model to learn and discern the underlying patterns within the data distribution. The boxplot indicates that the depth distribution has a broader interquartile range (IQR) and longer whiskers, suggesting greater complexity and noise. Meanwhile, the IQR for dimension height, depth error, and height error is narrower, which illustrates that these distributions are relatively compact with plain patterns.
Different geometry errors.  (1) Depth error $Z_{err}$ and dimension height error $H_{err}$ : Both $Z_{err}$ and $H_{err}$ have a narrow IQR and short whiskers, indicating that they share similar statistical properties. The primary distinction is that $Z_{err}$ distribution is symmetrical with a median close to zero, while $H_{err}$ distribution is skewed with a negative median. This skewness may increase the complexity of the training process. According to Eq. 3 and Eq. 5, we can express $H_{err}$ in terms of $Z_{err}$ and $h_{bbox}$:
${H_{err}}=\frac{1}{f}\cdot(Z-{Z_{geo}})\cdot{h_{bbox}}=\frac{{{Z_{err}}\cdot{h_{bbox}}}}{f}$ (7)
As illustrated in Fig. 4(b), when the object rotates around the center point, $Z_{err}$ will fluctuate in tandem with $h_{bbox}$, leading to an amplification effect on the distribution of $H_{err}$. Similar to the squaring operation in the chi-square distribution, the multiplication in Eq. 7 introduces nonlinear transformation, resulting in a skewed data distribution. (2) Bounding box height error $h_{err}$ : Compared with other geometry errors, $h_{err}$ is located in the denominator of the modified Eq. 6. Its gradient calculation formula can be expressed as:
$\frac{{\partial{L_{depth}}}}{{\partial{h_{err}}}}=\frac{{\partial{L_{depth}}}}{{\partial Z}}\cdot\frac{{f\cdot H}}{{{{({h_{bbox}}-{h_{err}})}^{2}}}}$ (8)
During the backpropagation process, the gradient calculations on $h_{err}$ are non-linear and unstable, thus increasing the difficulty of model training and convergence.
In summary, we obtain $Z_{geo}$ using $h_{bbox}$ instead of $h_{c}$ and select $Z_{err}$ to compensate the projection formula. These measures can greatly optimize data distribution and reduce the model’s learning costs. The difficulty of predicting depth-related attributes is ranked as follows:
$Z>{h_{err}}>{H_{err}}>{Z_{err}}\approx H>>{h_{bbox}}$ (9)
3.6 Loss Function
The training loss consists of four distinct components: region loss $L_{region}$, depth map loss $L_{dmap}$, 2D loss $L_{2D}$, and 3D loss $L_{3D}$. Specifically, $L_{2D}$ contains the losses of object category, projected center point, and bounding box, while $L_{3D}$ includes the losses of orientation, 3D size, and central depth. We denote the weights of these losses from ${\lambda_{1}}$ to ${\lambda_{6}}$. The overall loss of MonoDGP is formulated as:
${L_{overall}}={L_{2D}}+{L_{3D}}+{\lambda_{7}}{L_{dmap}}+{\lambda_{8}}\sum\limits_{i=0}^{4}{L_{region}^{i}}$ (10)
4 Experiments
4.1 Settings
Dataset.  We select the widely used KITTI benchmark [11] to evaluate our proposed model. This 3D object detection benchmark consists of 7481 training images and 7518 testing images. It has three difficulty levels (Easy, Moderate, and Hard) and three classes (Car, Pedestrian, and Cyclist). According to the prior work [7], we divide the 7481 training images into a training set with 3712 images and a validation set with 3769 images for ablation study.
Evaluation Metrics.  We present the detection outcomes categorized by three difficulty levels: easy, moderate, and hard. Consistent with prior works, our evaluation relies on average precision (AP) metrics for both 3D bounding boxes and bird’s-eye view, referred to as $AP_{3D}$ and $AP_{BEV}$, which are calculated at 40 recall positions following the established protocol [39]. The benchmark ranks all methods based on the moderate $AP_{3D}$ metrics of "Car" category. An IoU threshold of 0.7 is also applied during evaluation.
Implementation Details.  We take ResNet50 [12] as our backbone. In regard to Transformer, we adopt 8 heads in each attention module and set the number of queries 50 with 4 sampling points in the multi-scale deformable attention. The weights of losses are set as {5, 2, 10, 1, 1, 1, 1, 1} for ${\lambda_{1}}$ to ${\lambda_{8}}$. On a single RTX 3090 GPU, using Mixup3D, we train MonoDGP for 250 epochs with a batch size of 8 and a learning rate of 2e-4. We utilize AdamW [25] optimizer with weight decay and decrease learning rate by a factor of 0.5 at epochs 85, 125, 165, and 225. To accelerate the model’s convergence, we adopt group-wise one-to-many assignment [6] with 11 groups. During inference, we filter out queries with category confidence lower than 0.2.
4.2 Main Results
To present the performance of our proposed method, we provide quantitative results on the KITTI test and val splits.
As shown in Tab. 1, we have compared the most advanced monocular 3D object detection methods in recent years. Without any additional data, MonoDGP comprehensively surpasses the baseline MonoDETR [49], and achieves competitive results in most metrics. On test set and val set, compared to all existing methods, our performance for $AP_{3D}$ and $AP_{BEV}$ under three-level difficulties improves by -0.14%, +1.05%, +0.65%, +0.80%, +1.55%, +1.18%, +1.54%, +1.25%, +1.33%, +1.92%, +1.73%, +1.87%, respectively. Except that easy $AP_{BEV}$ on test set achieves the second-best lower than OccupancyM3D [29], other metrics all reach the SOTA performance.
It should be noted that we utilize early stopping to select the best checkpoint on val set, while we directly select the last few checkpoints and submit our results on test set. Considering the instability of the DETR-based model, on val set, we train MonoDGP 5 times and take the median of moderate $AP_{3D}$ scores as the final results. The rest of the ablation experiments are conducted in the same manner.
4.3 Ablation Study
Table 2: The design of region segmentation head. ‘w/o’ denotes enhancing visual and depth features without segment embeddings.
Segment Embeddings Threshold Val, $AP_{3D}$
Easy Mod. Hard
0.3 29.77 21.51 18.35
0.5 30.57 21.64 18.74
0.7 30.35 21.54 18.55
w/o 29.63 21.43 18.21
Region segmentation head.  For the multi-scale target region probabilities predicted by RSH, we set distinct thresholds to produce segment embeddings, which serve as input tokens for the depth encoder. The results are presented in Tab. 2. The introduction of segment embeddings strengths encoders’ ability to comprehend the contextual relationships among image pixels. Experimental findings further highlight the critical role of segment embeddings. Moreover, the threshold settings also influence 3D evaluation metrics. A threshold of 0.5 achieves superior results compared to thresholds of 0.3 and 0.7.
Table 3: Ablation study of different depth prediction modes on KITTI val set. The weighted fusion mode integrates the results from the depth map, direct depth, and geometric depth. "GD" denotes the geometric depth calculated by the formula $Z_{geo}=f*H_{3D}/h_{bbox}$.
Depth Predition Mode Val, $AP_{3D}$
Easy Mod. Hard
Depth Map 24.33 18.87 15.31
Direct Depth 26.60 19.23 16.47
Geometric Depth 9.15 7.99 5.52
Weighted Fusion 30.57 21.64 18.74
GD + Bbox Height Error 28.45 21.18 17.91
GD + Dim Height Error 30.46 22.09 18.85
GD + Depth Error 30.76 22.34 19.01
Depth prediction mode.  Based on the same network architecture, we predict depth in different modes, and the ablation results are expressed in Tab. 3. The findings indicate that individual depth estimations, whether using direct depth, depth map, or geometric depth, generally perform poorly, especially for geometric depth alone. As only perspective transformation on the visual surface considered, the moderate $AP_{3D}$ for geometric depth is just 7.99% (all $AP_{3D}$s mentioned below default to the moderate level), which is significantly lower than other prediction modes. By fusing the direct depth, depth map and geometric depth, which is adopted by MonoDETR, the $AP_{3D}$ improves to 21.64%. When geometry errors are added to compensate the projection formula, it can be observed that the network’s performance achieves a substantial increase. The depth error, with an $AP_{3D}$ of 22.34%, wins the best performance among all depth prediction modes. The dimension height error, with 22.09% in $AP_{3D}$, achieves the second-best performance due to its skewed distribution. The bounding box height error only acquires 21.18% in $AP_{3D}$, obviously inferior to other geometry errors. This discrepancy can be attributed to the nonlinearity of the gradient backpropagation calculation.
Table 4: The main ablation study.
Mixup3D 2D Query Region Segment Geometric Depth Error Val, $AP_{3D}$
Easy Mod. Hard
- - - - 26.56 20.22 17.21
- - - 28.17 20.79 17.52
- - 29.82 21.38 18.04
- 30.57 21.64 18.74
30.76 22.34 19.01
Improvement + 4.20 +2.12 +1.80
Overall performance.  We explore the impact of each improvement technique on the baseline. The main ablation study is shown in Tab. 4. The MonoDGP employs four key techniques: mixup3D augmentation, decoupling and initialization of 2D object queries, region enhancement with segment embeddings, and depth error prediction based on geometric depth. Compared to the baseline, the $AP_{3D}$s under three-level difficulties gain an increase by +4.20%, +2.12% and +1.80%, respectively. The experiments validate the effectiveness of each technique utilized in MonoDGP.
5 Conclusion
In this paper, we propose a novel transformer-based monocular method called MonoDGP, which adopts geometry errors to correct the projection formula. We also comprehensively analyze the properties and effectiveness of geometry errors, from the perspectives of marginal feature extraction, data distribution, and gradient propagation. Our error prediction is expected to become a better alternative to multi-depth prediction, without complex implementation of various depth branches. Furthermore, MonoDGP introduces a 2D visual decoder for query initialization and constructs a region segmentation head for feature enhancement. These improvements have achieved superior performance on the KITTI benchmark. Additionally, we can extend our approach to dense depth map prediction for the target region, by only considering errors between geometric depth and surface points of the objects.